# Topological and statistical signatures for quantifying ambiguity in semantic search

### v1

@@ 6 Conclusion and Next Steps
We have designed an experiment to identify and isolate topological signatures of ambiguity in a sentence, itself defined as a generalisation of words polysemy. We identified at least two signals derived from homologies $H_{0}$ and $H_{1}$ that exhibit different behaviour in ambiguous and unambiguous conditions. In addition to the discriminative utility of those signals they inform us of the geometric and topologic arrangement of query sentences neighbourhood which can be used as another probing tool of Large Language Models representation and structuring of information. As those signature seem independent of the language model used in our experiment, further investigations will explore how this relationship holds across a broader range of embedding dimensions and models, aiming to refine our understanding of semantic domain analysis in textual data. It is crucial to understand this in order to properly evaluate uncertainty of similarity metrics widely used in semantic search pipelines such as RAG (Gao et al. (2024)).
@@ Acknowledgments
Special thanks to: Javier Makmuri, Carlos Tomei, and Rob Murray-Rust for helpful discussions and varying levels of technical and emotional support. Special thanks to kids and spouses too (writing time is time away from family).
Disclosure. All opinions are our own and don’t represent that of BlackRock. Results here are not meant for investment purposes, and authors cannot be held liable to any misuse of the concepts proposed here.
Appendices
@@ Appendix A Geometric and Topological Foundations in Query Analysis
This appendix delves into the geometric and topological underpinnings essential to our analysis, particularly focusing on the application of homology theories to discern patterns within data clusters, cycles, and cavities. Our investigation employs the giotto-tda library, a versatile tool for Topological Data Analysis (TDA), enabling us to explore complex structures within high-dimensional data spaces. The rationale behind utilizing giotto-tda lies in its robust computational framework for extracting topological features, pivotal for understanding ambiguities in query datasets.
@@ A.1 Understanding Homologies
  • 3-Homology, our study’s top homology, identifies complex 3D simplices, necessitating at least five points for construction.
A.2 Application of giotto-tda
The giotto-tda library is instrumental in computing these topological features. By employing its Vietoris-Rips complex generator, we analyze the renormalized embeddings of our dataset to produce persistence diagrams. The Vietoris-Rips complex is a simplicial complex that is pivotal for understanding the shape of data. It is constructed by connecting points that are within a certain distance from each other, thus capturing the underlying topological structure of the data at different scales.
These diagrams, in turn, facilitate the examination of homological features such as the birth of homologies, indicative of the scale at which points in a point cloud start to connect, reflecting the overall density and distribution of points.
A.3 Derivation of 1-Wasserstein Norm
The Wasserstein distance is typically used for comparing dissimilarities between distributions. More specifically it has a use case to compare topological spaces from point clouds by the means of quantifying similarities between persistence diagrams (Kerber et al. [2016], Wasserman [2016]) for each homology orders separately. As an example for two persistent diagrams A and B of homology H0, the general formula of the Wasserstein distance in the "giotto-tda" (Tauzin et al. [2021]) package is written as:
@@ A.3 Derivation of 1-Wasserstein Norm
[…]
@@
A.4 Renormalization
For a given scaling factor $\varepsilon$, the distributions of nearest neighbours are spanning over a different fraction of the top-$k$ retrieval results (with $k=50$) between unambiguous and ambiguous configuration as shown in figure 6. As much as this density of neighbours is already indicative of whether a query aggregates different topics we want to compare ambiguous from unambiguous queries with the same settings and therefore for each value of $\varepsilon$ we limit ourselves to a number of nearest neighbours bounded by $\min(\max(\mathcal{N}_{ambiguous}),\max(\mathcal{N}_{unambiguous}))$ with $\mathcal{N}$ being the set of neighborhood sizes.
In addition those distributions are not uniform because we picked an arbitrary sample of documents to build $\mathcal{D}_{BLK}$ which is why we correct the persistent homology features distributions by weights corresponding to:
$w=1/\rho_{queries}(N_{NearestNeigbours},\varepsilon)$
where $\rho_{queries}(N_{NearestNeigbours},\varepsilon)$ is the density of queries function of the neighbourhood size and the scaling factor $\varepsilon$
Figure 6: a) Empirical distribution of queries neighbourhood size with scaling factor $\varepsilon=0.3$ for ambiguous and ambiguous configurations in blue and red respectively. b) same distribution for a scaling factor $\varepsilon=4$
A.5 Pseudo-code for Calculating Homology features
@@
1:Input: Renormalized embeddings $\bm{v_{E}}$
2:Output: Homology 0 and 1 metrics, $W_{1}(H_{0})$ and $LT_{max}(H_{1})$
3:procedure CalculateHomologyMetrics($\bm{v_{E}}$)
4:     Initialize Vietoris-Rips complex with the renormalized embeddings $\bm{v_{E}}$
5:     Fit the complex to the embeddings
6:     Extract the first persistence diagram
7:     Extract homology (birth,death) ordered pairs from the diagram
8:     Compute $W_{1}(H_{0})$ using Wasserstein distance
9:     Compute $LT_{max}(H_{1})$ using the lifetime of the longest feature
10:end procedure
@@
Incorporating these insights enhances our understanding of the dataset’s structure, offering a nuanced view of how topic diversity impacts the geometric and topological features we observe. Through this detailed analysis, we aim to provide a comprehensive framework for assessing query ambiguity, leveraging the sophisticated tools offered by giotto-tda to navigate the complexities of high-dimensional data spaces.
Appendix B Geometric intuition of topological features
The 1-Wasserstein norm of connected components (i.e.: $W_{1}(H_{0})$) should carry some information about the anisotropy and the local density of the projected query neighbourhood on the hypersphere $S^{n}$. Nevertheless it is non-trivial to directly infer geometric information from that quantity. In order to build our intuition we considered a toy model where we define:
  • The dimensionality of the space $D$
  • The number of neighbours $N_{\text{neigh}}$ we want to represent.
  • Their equipartition in $N_{\text{clust}}$ clusters.
Together, $N_{\text{neigh}}$ and $N_{\text{clust}}$ provide a control on the density of the neighbourhood. In addition we define a scaling factor $\alpha$ that biases the sampling of a neighbour vector along a specific dimension $i$. All neighbours from the same cluster are sampled with the bias along the same dimension. By doing so we can control the anisotropy of the neighbourhood distribution. $W_{1}(H_{0})$ can therefore be calculated for each quadruplet of parameters $(D,N_{\text{neigh}},N_{\text{clust}},\alpha)$ by the algorithm:
Algorithm 3 Anisotropic $W_{1}$ sampling in the space $S^{n}$
1:Input: $D,N_{\text{neigh}},N_{\text{clust}},\alpha$
2:Output: $W_{1}(H_{0})$
3:procedure SampledWasserstein($D,N_{\text{neigh}},N_{\text{clust}},\alpha$)
4:     For $i$ in $range(N_{\text{clust}})$:
5:           Define the covariance matrix $\Sigma$ as the Identity matrix $I$
6:           Substitute $\Sigma_{ii}=1$ for $\Sigma_{ii}=\alpha$
7:           Draw $N_{\text{neigh}}/N_{\text{clust}}$ sample vectors $\vec{v}_{k}=(x_{0},\dots,x_{n-1})$ with $x_{j=i}$ sampled from $\mathcal{N}(\vec{0},\Sigma)$.
8:           Renormalize each vector $\vec{v}^{{}^{\prime}}_{k}=\frac{\vec{v}_{k}}{||\vec{v}_{k}||}$
9:     Compute Vietoris-Rips persistence diagram for the set $\{\vec{v}^{{}^{\prime}}_{k}\}_{k=0}^{N_{\text{neigh}}}\rightarrow H_{0}$
10:     Compute $W_{1}(H_{0})$
11:end procedure
and we can observe its behaviour in different neighbourhood configurations in figures 7 and 8.
Figure 7: Evolution of the norm $W_{1}(H_{0})$ with scaling coefficient $\alpha$ for $D=768$, $N_{\text{clust}}=3$ and variable $N_{\text{neigh}}$. On top of the main figure are example of biased distributions along the 3 axes in $S^{2}$ with different $\alpha$ values (blue circles). 3 clusters are sampled along those dimensions (red crosses)
Figure 8: Evolution of the norm $W_{1}(H_{0})$ with scaling coefficient $\alpha$ for $D=768$ and variable $N_{\text{neigh}}$ and $N_{\text{clust}}$.
Note that biasing the distribution of each cluster along an independent dimension of the space $S^{n}$ corresponds to a boundary case of a real context where every dimension encodes mainly a single semantic. It also means that the cosine similarity across clusters tends to $\sim 0$ as $\alpha$ increases. The configuration $(N_{\text{neigh}}=2,N_{\text{clust}}=1)$ is trivial to analyze as $W_{1}(H_{0})$ only represents the scaled distance between the two points, which drops as those two points are sampled with a stronger orientation bias accounting only for the anisotropy of the neighborhood. The same conclusions can be drawn for all configurations with $N_{\text{clust}}=1$.
For all other configurations, we can distinguish three different regions in figure 8:
  • Region I: High $W_{1}(H_{0})$ values are achieved irrespective of the number of clusters ($N_{\text{clust}}$) and the number of neighbors ($N_{\text{neigh}}$). This corresponds to a more uniform sampling (strictly uniform in the case of $\alpha=1$). In a real context, this behavior would be expected when data is distributed uniformly across all dimensions in high-dimensional spaces.
  • Region II: This marks a transition regime with a clear drop in the norm. The separation between clusters becomes larger, but intra-cluster distances also shrink (indicating higher cluster density), which significantly impacts $W_{1}(H_{0})$. The neighborhood becomes more structured and anisotropic.
  • Region III: This region exhibits a relative stabilization of $W_{1}(H_{0})$ but with high variance between different $(N_{\text{neigh}},N_{\text{clust}})$ configurations. Increasing the number of clusters raises the norm value, while a higher density within each cluster decreases it. Therefore, anisotropy is mitigated by the increasing number of clusters.
Appendix C Embedding Models
The three models we have used, SBERT-multiqa v6, SBERT-msmarco v4, and GTR-T5-XL, vary in architecture and application but all three have same output dimensionality equal to 768. The SBERT-multiqa (Reimers and Gurevych [2019], Huggingface [2022]) is a pretrained multi-qa-mpnet-base-cos-v1 model designed for efficient similarity search in multi-domain question-answering. The SBERT-msmarco (Reimers and Gurevych [2019], Huggingface [2021]) is also sentence-bert model fine-tuned on the MS MARCO dataset (Bajaj et al. [2018]), targeting high relevance in retrieval tasks. Meanwhile, the GTR-T5-XL (Ni et al. [2021]) leverages a large T5 transformer, optimized for general textual retrieval across a wider range of datasets and tasks.
Appendix D Hyperparameters of computational proxy for semantic domains
In order to define the semantic domains (or topics) we performed dimensionality reduction followed by clustering using UMAP and HDBSCAN. We chose UMAP because it can retain both local or global topological structure. We performed the reduction down to 2 dimensions and optimised it on the parameter n__\__neighbors which corresponds to the number of neighbours the model will consider. Intuitively, low values favor local structure projection and high values favor global one. We chose to automate the search of optimal n__\__neighbors for each document independently by scanning a range $n_{neighbors}\in(2,50]$ and considering the maximum rate of pairwise distance $\delta_{ij}$ change:
$n^{*}_{\text{neighbors}}=\arg\max_{n_{\text{neighbors}}}\left(\frac{d^{2}\left(\sum_{i,j;i\neq j}\delta_{ij}(n_{\text{neighbors}})-\delta_{ij}(2)\right)}{dn_{\text{neighbors}}^{2}}\right)$ (7)
Notation Description
$n_{\text{neighbors}}$ The number of neighbors considered by the UMAP model.
$\delta_{ij}$ Pairwise distance between points $i$ and $j$ in the reduced dimensional space.
$\delta_{ij}(n_{\text{neighbors}})$ Pairwise distance calculated using a specific $n_{\text{neighbors}}$.
$\delta_{ij}(2)$ Pairwise distance when $n_{\text{neighbors}}=2$, used as a reference point.
$\arg\max$ Indicates that we are searching for the value of $n_{\text{neighbors}}$ that maximizes the expression.
$\frac{d^{2}(\cdot)}{dn_{\text{neighbors}}^{2}}$ Second derivative with respect to $n_{\text{neighbors}}$, capturing the rate of change.
Table 3: Summary of Notation
This approach empirically favours relatively low parameter value (n__\__neighbors $\simeq$ 10) and ensure that the local topological structure of the embeddings is preserved in the reduction process while scaling easily the experiment to an arbitrary number of documents. The "min__\__dist" parameter is left at its default value of 0.1. A value too low ($\simeq$0.0) would result in collapsing of similar embeddings in single positions in 2D while a high value would lead to a strong spread of data points which would make the optimisation process in 7 impossible.
[…]

### v2

@@ 6 Conclusion and Next Steps
In this article we have defined a formal definition of ambiguity in the context of text retrieval as the intersection of topics contained in the query and its retrieved answers candidates. We evidenced two main regimes: polysemic and multi-factual. From there, using the linear representation hypothesis to encode the topics, we constructed synthetic datasets that match the different ambiguity regimes. We evidenced that, similarly to previous works on words polysemy, the different regimes of ambiguity exhibit clear and separable persistent homology $H_{0}$ and $H_{1}$ fingerprints. Finally we validated those findings with real-world dataset experiment where we isolated the same topological signatures of ambiguity in a sentence. In addition to the discriminative utility of those signals they inform us of the geometric and topological arrangement of query sentences neighborhood which can be used as another probing tool of Large Language Models representation and structuring of information. As those signature seem independent of the language model, further investigations will explore how this relationship holds across a broader range of embedding dimensions and models, aiming to refine our understanding of semantic domain analysis in textual data. It is crucial to understand this in order to properly evaluate uncertainty of similarity metrics widely used in semantic search pipelines such as RAG (Gao et al. (2024)).
@@ Acknowledgments
Special thanks to: Paul Gaskell and Firuza Mamedova for helpful discussions and varying levels of technical support.
Disclosure. All opinions are our own and don’t represent that of the employers we work and/or worked for. Results here are not meant for investment purposes, and authors cannot be held liable to any misuse of the concepts proposed here.
@@ Appendix A Geometric and Topological Foundations in Query Analysis
This appendix delves into the geometric and topological underpinnings essential to our analysis, particularly focusing on the application of homology theories to discern patterns within data clusters, cycles, and cavities. Our investigation employs the Ripser library enabling us to explore complex structures within high-dimensional data spaces. The rationale behind utilizing Ripser lies in its robust computational framework for extracting topological features, pivotal for understanding ambiguities in query datasets.
@@ A.1 Understanding Homologies
  • 3-Homology, our study’s top homology, identifies complex 3D simplexes, necessitating at least five points for construction.
A.2 Derivation of 1-Wasserstein Norm
The Wasserstein distance is typically used for comparing dissimilarities between distributions. More specifically it has a use case to compare topological spaces from point clouds by the means of quantifying similarities between persistence diagrams (Kerber et al. (2016); Wasserman (2016)) for each homology orders separately. As an example for two persistent diagrams A and B of homology $H_{0}$, the general formula of the Wasserstein distance in the ”giotto-tda” (Tauzin et al. (2021)) package is written as:
@@ A.2 Derivation of 1-Wasserstein Norm
where $\Delta$ corresponds to the virtual set of diagonal points in the diagram that ensures that there is a full bijection between A and B, and $\|.\|_{\infty}$ accounts for $\max(|x|,|y|)$ with $(x,y)\in\mathbb{R}^{2}$ the coordinate of a point in the diagram. In our experiments we use $W_{1}$, homology $H_{0}$ so $\max(|x|,|y|)\rightarrow|y|$ as all components lie on the vertical axis $(x=0,y)$ and instead of comparing two diagrams A and B we will calculate the norm for diagram A only which means that we replace B by the empty set $\{\emptyset\}$. Finally we can discard the infimum condition as the orthogonal projection of any point on $(x=0,y)$ onto the diagonal axis $(x,y=x)$ is the function that minimizes the distances. Therefore the formula transforms into:
@@
A.3 Pseudo-code for Calculating Homology features
Input: Re-normalized embeddings $\bm{v_{E}}$
Output: Homology 0 and 1 metrics, $W_{1}(H_{0})$ and $LT_{\text{max}}(H_{1})$
Procedure CalculateHomologyMetrics($\bm{v_{E}}$):
   Initialize Vietoris-Rips complex with the re-normalized embeddings $\bm{v_{E}}$;
   Fit the complex to the embeddings;
   Extract the first persistence diagram;
   Extract homology (birth, death) ordered pairs from the diagram;
   Compute $W_{1}(H_{0})$ using Wasserstein distance;
   Compute $LT_{\text{max}}(H_{1})$ using the lifetime of the longest feature;
@@ A.3 Pseudo-code for Calculating Homology features
Incorporating these insights enhances our understanding of the dataset’s structure, offering a nuanced view of how topic diversity impacts the geometric and topological features we observe. Through this detailed analysis, we aim to provide a comprehensive framework for assessing query ambiguity, leveraging the sophisticated tools offered by persistent homologies to navigate the complexities of high-dimensional data spaces.
Appendix B ”Ab-Initio” Simulated Persistent Homology
This appendix describes an simulation framework for analyzing persistent homology features in hierarchical cluster structures generated through orthogonal feature-based embeddings. We investigate three scenarios that model different query-corpus relationships commonly encountered in document retrieval and semantic search systems: (1) parent queries to children corpora, (2) child queries to parent corpora, and (3) parent queries to children with variable correlations. The framework systematically varies noise levels $\sigma_{\text{noise}}$, dimension ratios ($n_{\text{parent}}/n_{\text{child}}$), neighborhood locality (scaling factor $\varepsilon$) and feature overlap to understand how topological features reflect semantic relationships.
B.1 Data Generation: Topic Vocabulary-Based Clustering
We generate synthetic hierarchical clusters using a vocabulary of orthogonal topic vectors. Each datapoint is constructed by sampling and summing topics from this vocabulary, then adding gaussian noise. This models document embeddings where semantic concepts (topics) are combined to create complex representations.
B.1.1 Orthogonal Topic Vocabulary Generation
The topics vocabulary is created using QR decomposition to ensure approximate orthogonality:
Input: NN: total number of topics, $D$: embedding dimension
Output: $\mathcal{F}=\{\mathbf{f}_{0},\ldots,\mathbf{f}_{N-1}\}$ where $\mathbf{f}_{i}\in\mathbb{R}^{D}$
$M\leftarrow$ random matrix $\in\mathbb{R}^{D\times n_{\text{topics}}}$ from $\mathcal{N}(0,1)$;
Compute QR decomposition: $M=QR$ where $Q\in\mathbb{R}^{D\times n_{\text{features}}}$ is orthonormal;
for $i=0$ to $n_{\text{topics}}-1$ do
 $\mathbf{v}_{i}\leftarrow Q[:,i]$ // Extract $i$-th orthonormal column
 $\text{threshold}\leftarrow\text{median}(\mathbf{v}_{i})$;
 $\mathbf{b}_{i}\leftarrow(\mathbf{v}_{i}\geq\text{threshold})$ // Binary mask: 1 if above median, 0 otherwise
 $\mathbf{f}_{i}\leftarrow\mathbf{b}_{i}/\|\mathbf{b}_{i}\|_{2}$
 $\mathcal{F}[i]\leftarrow\mathbf{f}_{i}$;
end for
return $\mathcal{F}$
Algorithm 2 Generate Orthogonal Topics Vocabulary
  • Orthogonality: Topics from the vocabulary are approximately orthogonal due to QR decomposition
  • Arbitrary basis: Topics are distributed across all $D$ dimensions of the embedding space reflecting real case where there is no trivial alignment between the embedding space basis and the basis defined by the topics.
B.1.2 Datapoint Generation with Feature Sampling
We construct a simple 2-level hierarchy for controlled experimentation:
  • Parent level: Clusters with $n_{\text{parent}}$ features (depth 0)
  • Children level: Clusters with $n_{\text{child}}<n_{\text{parent}}$ features (depth 1)
Each datapoint is generated by sampling a subset of features from the feature vocabulary and adding variable gaussian noise:
Input: $k$, $\mathcal{I}_{\text{parent}}=\{i_{1},\ldots,i_{n_{\text{parent}}}\}$ or $\{\emptyset\}$, $\mathcal{F}$, $n_{\text{points}}$, $\sigma_{\text{noise}}$
Output: $X\in\mathbb{R}^{n_{\text{points}}\times D}$
Initialize $X\in\mathbb{R}^{n_{\text{points}}\times D}$;
for $j=1$ to $n_{\text{points}}$ do
 if $\mathcal{I}_{\text{parent}}=\{\emptyset\}$ then
      Sample $k$ topics: $\{\mathbf{f}_{j,1},\ldots,\mathbf{f}_{j,k}\}$ from $\mathcal{F}$ without replacement;
   end if
 else
      Sample $k$ topics: $\{\mathbf{f}_{j,1},\ldots,\mathbf{f}_{j,k}\}$ from $\{\mathcal{F}[i]:i\in\mathcal{I}_{\text{parent}}\}$ without replacement;
   end if
  Sum topics: $\mathbf{x}_{j}\leftarrow\sum_{\ell=1}^{k}\mathbf{f}_{j,\ell}$;
 if $\sigma_{\text{noise}}>0$ then
      Sample noise: $\bm{\epsilon}_{j}\sim\mathcal{N}(0,\sigma_{\text{noise}})$
     Add noise: $\mathbf{x}_{j}\leftarrow\mathbf{x}_{j}+\bm{\epsilon}_{j}$;
   end if
  Normalize: $\mathbf{x}_{j}\leftarrow\mathbf{x}_{j}/\|\mathbf{x}_{j}\|_{2}$;
 $X[j,:]\leftarrow\mathbf{x}_{j}$;
end for
return $X$
Algorithm 3 Generate Datapoints from Topics
  • Additive composition: Datapoints are weighted sums of sampled topics
  • Controlled hierarchy: Parent-child relationships with explicit topics count control
  • Topic overlap testing: Can generate orthogonal or overlapping parent-child pairs
B.2 Robustness of the results
Different simulation conditions have been tested and exhibit consistent persistent homology pattern.
Figure 5: Top Left) Simulated $W_{1}(H_{0})$ for scenario 1 (red) and scenario 2 (blue). Top Right) Simulated $LT_{max}(H_{1})$ for scenario 1 (red) and scenario 2 (blue). Bottom Left) Simulated $W_{1}(H_{0})$ for scenario 1 (red) and scenario 2 (blue) as a function of the relative neighborhood scaling factor $\varepsilon$. Bottom Right) Simulated $LT_{max}(H_{1})$ for scenario 1 (red) and scenario 2 (blue) as a function of the relative neighborhood scaling factor $\varepsilon$. Varied simulation conditions: $\{(D=64,N=16,n_{\text{parent}}=12),(D=128,N=32,n_{\text{parent}}=16),(D=256,N=64,n_{\text{parent}}=32)\}$, Fixed simulation conditions: $\varepsilon=0.4,\sigma_{\text{noise}}=0.1$
We also performed an additional test of robustness by applying dimensionality reduction to our embedding vectors. According to the Johnson-Lindenstrauss lemma, a random projection to a lower dimension space $D^{\prime}<D$ preserves the distances between the datapoints and we expect the homology signal to be stable under that transformation. We report here that our simulation homology signal is invariant with respect to such a random projection (figure 6)which reinforce our claim that it depends on the relative topics encoding of datapoints and not the absolute dimensionality of the embedding space.
Figure 6: Projecting query and corpus embeddings from 256 dimensions to $\{224,192,128,64\}$ dimensions with random projection matrix. Starting Simulation conditions: ($D=256,N=64,n_{\text{parent}}=32,\varepsilon=0.4,\sigma_{\text{noise}}=0.01$)